In a triangle the measure of the largest angle is 3 times the measure of the smallest angle. The measure of the remaining angle is twice the measure of the smallest angle. What is the measure of the longest angle?
Answer by texttutoring(324)

Let x = first angle
Let y = second angle
Let z = 3rd angle

We know that in a triangle, the 3 angles have to add up to 180 degrees:

x+y+z = 180

We also know that:
z = 3x
y = 2x

Plug this information into the other equation:

x + y + z = 180
x + 2x + 3x = 180
6x = 180
x = 30

This means y=2(30)=60 and z=3(30)=90

x=30
y=60
z=90